OpenStudy (meowlover17):

What is the Greatest Common Factor of x6 and x4?

OpenStudy (meowlover17):

x^6 and x^4

OpenStudy (meowlover17):

x x4 x6 x10

OpenStudy (meowlover17):

I think its X

OpenStudy (campbell_st):

well think of it this way \[x^6 = x^4\times x^2\] this may help you identify the GCF

OpenStudy (meowlover17):

well based off of my answer choices i think its X

OpenStudy (meowlover17):

but it would be x^2

Nnesha (nnesha):

variable with the lowest degree always going to be GCF

Nnesha (nnesha):

that's how i remember it ;P

OpenStudy (meowlover17):

So is it x?

jimthompson5910 (jim_thompson5910):